# iRAG: An Incremental Retrieval Augmented Generation System for Videos

### v1

@@ Abstract
Retrieval augmented generation (RAG) systems combine the strengths of language generation and information retrieval to power many real-world applications like chatbots. Use of RAG for combined understanding of multimodal data such as text, images and videos is appealing but two critical limitations exist: one-time, upfront capture of all content in large multimodal data as text descriptions entails high processing times, and not all information in the rich multimodal data is typically in the text descriptions. Since the user queries are not known apriori, developing a system for multimodal to text conversion and interactive querying of multimodal data is challenging.
@@
To address these limitations, we propose iRAG, which augments RAG with a novel incremental workflow to enable interactive querying of large corpus of multimodal data. Unlike traditional RAG, iRAG quickly indexes large repositories of multimodal data, and in the incremental workflow, it uses the index to opportunistically extract more details from select portions of the multimodal data to retrieve context relevant to an interactive user query. Such an incremental workflow avoids long multimodal to text conversion times, overcomes information loss issues by doing on-demand query-specific extraction of details in multimodal data, and ensures high quality of responses to interactive user queries that are often not known apriori. To the best of our knowledge, iRAG is the first system to augment RAG with an incremental workflow to support efficient interactive querying of large, real-world multimodal data. Experimental results on real-world long videos demonstrate 23x to 25x faster video to text ingestion, while ensuring that quality of responses to interactive user queries is comparable to responses from a traditional RAG where all video data is converted to text upfront before any querying.
Index Terms:
@@
Video Analytics, Retrieval Augmented Generation (RAG), Large Language Model, Generative AI, Multimodal Data, Natural Language Processing
I Introduction
@@
Many applications routinely collect and store videos for offline analysis. For example, surveillance systems ensure public safety by monitoring public spaces [1], airports, transportation hubs, and critical infrastructure using video cameras. They also store video feeds for offline analysis and investigations. There is also an increasing usage of video analytics for patient monitoring in hospitals and healthcare facilities. Analyzing video feeds from traffic cameras for traffic monitoring, congestion management [2], and incident detection also requires an understanding of information in videos.
@@
Large language models (LLMs) like ChatGPT [3] have demonstrated their proficiency in human-like text-based dialogues. Inspired by the success of LLMs in understanding large repositories of text, recent proposals attempt to understand information in videos by using LLMs [4, 5]. These proposals use vision AI models to analyze content in each video frame, and these models output the visual information in a frame as a text description. For example, an object detector AI model can identify objects in each frame as cars, trucks, bicycles, persons etc., and also output the location of these objects in a frame as human understandable text. Such AI models can process long videos and capture visual information as a text document, which is a concatenation of text descriptions from each frame. LLMs can leverage these text documents using advanced techniques like retrieval augmented generation (RAG) [6], and provide useful responses to interactive queries about the content in the videos.
@@
Prior proposals to understand video content through interactive queries have two major limitations. First, the time required to process entire long videos and generate text descriptions by using complex AI models can be prohibitive. For example, it could take more than a day for complex AI models to analyze visual information in a 24-hour surveillance video and generate text descriptions. This prevents the police from timely analysis of criminal incidents in the video. Second, not all visual information in the videos is captured in the text descriptions, even if we use state of the art complex vision AI models [7, 8, 9, 10]. Furthermore, the user queries are not known apriori, and it is difficult to determine which AI models should be used to convert visual content into text.
@@
To address the critical limitations of prior efforts to interactively query content in long videos, we propose iRAG. Unlike prior proposals [4, 5] that employ many AI models to extract textual information upfront, iRAG uses lightweight AI models to quickly prepare an index for the content in the long videos. During preprocessing of the long video, our primary goal is to extract enough information from the video to be able to index the video data. Such an index facilitates effective extraction and retrieval of more detailed information from select sections of the video in response to interactive user queries.
@@
While utilizing lightweight AI models reduces the indexing time, the text descriptions may not be as rich, and many user queries may not be answerable by an LLM due to lack of facts in the text descriptions. To tackle this challenge, we propose an incremental workflow. When the information extracted during indexing is inadequate to successfully respond to a user query, additional models are used on-demand to extract more details from selective parts of the long video. This strategy ensures that iRAG extracts necessary detailed textual information on-demand, addressing specific queries as they arise, rather than collecting all textual information upfront without knowing the specific details needed to answer arbitrary user queries.
@@
  1. 1.
    We propose a novel incremental RAG system called iRAG to quickly start interactive querying of large video repositories. Unlike traditional methods that employ many AI models up front to convert entire video information into text, iRAG makes query-aware selection of AI models as necessary. Compared to prior work, iRAG improves the preprocessing time by 23x to 25x, and uses incremental extraction to ensure comparable useful LLM responses.
  2. 2.
    iRAG consists of three novel components that work together to enable incremental RAG: (a) Given a user query, the Query Planner identifies sections of the video that should be analyzed in more detail. The Planner also identifies the AI models that should be used for detailed analysis. (b) The Indexer uses a novel AI-based re-ranking method to significantly reduce the portions of the long video that should be considered for detailed analysis. (c) Finally, the Extractor extracts detailed information, and updates the index to ensure query-relevant retrieval of context that is useful to elicit useful responses from an LLM.
II Motivation
Prior work [4, 5] has augmented user prompts with additional context from stored videos to guide an LLM to produce more contextually relevant and informed responses. These systems produce text descriptions of content in long videos upfront by using a variety of AI models, and they use traditional RAG systems to support interactive querying. We highlight two key limitations that arise when existing methods are applied to real-world videos and scenarios, and we propose a novel, incremental workflow to address them.
II-A Prior work to leverage context in video
Prior work [4, 5] first produces a textual description of information in the entire video. One or more AI models like object detection analyze every frame in the stored video and produce a textual description like a list of objects in the frame. Concatenation of descriptions for each frame results in a long document. Then, a traditional RAG system is used to augment the user prompt with query-specific context in the long document. The augmented prompt guides the LLM to produce more useful responses. We briefly review this process.
Preprocessing: As shown in Figure 1, a long video is divided into smaller clips. Each clip has a pre-determined number of frames. Each frame in a clip is analyzed by one or more vision AI models like object detection to produce a textual description. Text from all the frames in the video is concatenated into a long document. A variety of vision or generative AI models can be used to extract information from each video frame. However, video to text conversion can take a long time depending on the available computational resources, and complexity of the AI models.
Figure 1: Video to text conversion
@@ II-A Prior work to leverage context in video
RAG: The long text document is stored in a vector database. This makes it easy to retrieve portions of the document that are most relevant to a given user query. The long text document is chopped up into chunks where a chunk can be a paragraph, or a sentence or some pre-determined number of consecutive characters. Each chunk is embedded into a vector of pre-determined length by an embedding algorithm, and these vectors are all stored in a vector database. All this makes it easy to identify chunks in the document that are similar to a user query.
For example, given a user query, we can embed the query (by using the embedding model) into a vector. Then, we can compute the dot product of the query vector with each vector in the database, and easily determine the top $k$ vectors in the database that are most similar to the query vector. As shown in Figure 2, $k$ is a configuration parameter of the baseline RAG system.
Since we know the chunk that corresponds to each vector in the database, it is easy to retrieve the $k$ chunks in the long document that are most similar to the user query. The semantic search block in Figure 2 retrieves the relevant query-specific context from the long document. Then, user query is augmented with query-specific context, and the augmented prompt results in a more contextually accurate response from an LLM.
In this paper, we use the preprocessing and retrieval methods shown in Figures 1 and  2 as our baseline to evaluate our proposed new approach.
Figure 2: Conventional RAG
II-B Limitations of Prior work
@@
Long pre-processing time: Since the user queries are not known apriori, all available vision AI models are used to glean information in the stored video. Often, models that can capture more details and nuances in the video are often very computationally expensive. Preprocessing time matters. For example, crimes are time-sensitive, and the sooner law enforcement can analyze surveillance footage, the greater the likelihood of capturing critical evidence while it is still fresh. As we show later, preprocessing of a 24-hour surveillance video using prior work can easily take more than a day. This limits how quickly the context in the videos can be used to solve real-world problems, especially when the timely analysis of surveillance video is crucial for law enforcement (to gather actionable intelligence, identify suspects, or conduct thorough investigations) to solvd crimes and ensure public safety.
Figure 3 shows the preprocessing time bottleneck for querying and analyzing long videos from several traffic intersections in New York city [11]. We consider three videos that are 1, 7, or 30 days long. We use the DETR object detection model [7] and the GRiT dense captioning model [10] to extract text information from the videos. The average per-frame latency times are approximately 70 ms, and 1500 ms for DETR and GRiT, respectively, on a server equipped with an AMD Ryzen 5950X 16-Core processor and a NVIDIA GeForce RTX 3090 GPU. Although we process only one frame in every one-second video clip, the video-to-text generation process for for 1-day, 7-day, and 30-day videos requires approximately 1.57 days, 11 days, and 46.4 days, respectively. Such long video-to-text conversion times hinder development of practical systems to quickly query and understand long videos from real-world applications.
@@
Figure 3: Time to generate a long document from real-world videos
@@
Our new approach: Instead of executing all vision AI models upfront to extract textual information from entire videos, we propose an incremental RAG (iRAG) system that does query-aware, on-demand extraction of text data from select portions of the videos. Furthermore, iRAG automatically selects appropriate vision AI models that are necessary during the incremental workflow to answer the user query.
Such a capability is not present in prior methods [4, 5] that solely rely on long preprocessing times to convert video to text once before any interactive querying of the long videos.
III Incremental RAG
III-A Preprocessing in iRAG
Unlike prior approaches, the goal of preprocessing in iRAG is to quickly build an index for the long video. For example, a lightweight object detection model like DETR [7] can analyze each frame in the video and identify objects, their types, their location and size (bounding box). Such information forms a natural index to quickly identify the clips in the video that may have information relevant to a user query. Frames provide a natural timeline, with each frame containing information that can be searched to identify relevant clips for responding to a user query.
For example, to answer the query “Is there a FedEx truck in a long video?”, the index (i.e. the object detection information) can be used to quickly identify the relevant clips where a truck was detected. Subsequent detailed extraction of text information from these select clips (by using an AI model that can pick up text in images) can confirm if and where a FedEX truck is seen in the long video. We remark that a variety of indexes can be quickly constructed by using lightweight vision AI models. Also, the indexes do not have to be human-readable or understandable, and in the future, it is likely that the features used to construct the index are automatically generated by AI models. However, for ease of understanding of our new approach, we limit the discussion to the use of two popular lightweight models (DETR [7] and CLIP [12]) during preprocessing in iRAG. Such an index is human understandable, and it provides transparency and explainability of the query responses from iRAG.
III-B Query-Response
Given a user query, iRAG retrieves the relevant context from the information collected during preprocessing by using the conventional RAG approach. Then, we construct a suitable prompt for the LLM by using the following prompt template:
Prompt 1
@@ III-B Query-Response
We invoke the incremental flow when the LLM is unable to answer a user query. Next, we describe the iRAG system, and its key components.
III-C System Overview
Figure 4 shows an overview of the proposed iRAG system. The incremental flow (which is invoked only when the LLM is unable to answer a query) is shown using dotted lines.
There are three key components that support the query-aware, on-demand incremental flow: query planner, indexer and extractor.
Figure 4: System overview of iRAG. The dotted lines represent the additional workflow beyond a conventional RAG.
Our Query Planner retrieves query-aware relevant context from the information gathered in preprocessing, and recommends AI models that should be used to extract more details from selected clips in the long video. These additional details are necessary to satisfactorily answer the user query. Context retrieved by the Planner consists of two parts: the top-$N$ chunks in TextDB (containing information from the DETR model) that are most similar to the user query, and the top-$F$ frames in ImageDB (containing information from the CLIP model) that are most similar to the user query. Note that $N$ and $F$ are configuration parameters for the iRAG system.
Our Indexer further refines (and reduces) the context from the Planner to produce at most $k$ video clips that will be analyzed the Extractor. These video clips are the context clips. As shown in Figure 4, $k$ is also a configuration parameter for the iRAG system. Unlike the the interpretation of $k$ in the baseline system (where $k$ chunks of text context are used to construct the prompt for the LLM), iRAG interprets $k$ to be the upper limit on the number of context clips that will be sent to the Extractor. Since detailed extraction of information from the context clips dominates the query processing time, the choice of $k$ affects the quality of interactive querying experience. Furthermore, the configuration parameters $N$ and $F$ are typically an order or magnitude larger than $k$ (as described in the Experimental Results section IV).
Our Extractor extracts more details from the context clips by using AI models recommended by the Planner. The text and image vectors obtained from the extraction of context clips are added to the TextDB and ImageDB, respectively, to update the index. We then re-run the user query by retrieving updated context, and the likelihood of eliciting a useful response from the LLM increases.
The iRAG system can be configured to run the incremental flow in a loop for a preset maximum number of times, after which the system reports that the user query cannot be answered. In practice, a single iteration of the incremental flow is adequate to successfully answer most queries (as observed in the results reported in the Experimental Results section IV).
The response to a query consists of two parts: a textual response to the user query, and a set of video clips from the long video that support the textual response (refer to Figure 11 for an illustration). This allows the user to detect incorrect responses (or hallucinations) from the LLM. In the future, we can also construct automatic response verifiers that determine whether the human understandable textual response and video clips support each other.
III-D Query Planner
@@
Figure 5 shows the workflow in our planner. Given a query, the Planner computes a query embedding vector by using the same embedding algorithm that the RAG system used to embed preprocessing text chunks into vectors in TextDB, and retrieves the top-$N$ vectors in TextDB that are most similar to the query embedding vector. Chunks that correspond to the top-$N$ vectors are the top-$N$ chunks, which are the most relevant chunks for the user query.
Figure 5: Query planner workflow
@@
Improved context retrieval: To improve the quality of the retrieved context, we use an additional vector database ImageDB to store frame vectors. Lightweight AI models like CLIP embed visual information in each frame of the long video into a CLIP embedding vector, which is also the frame vector. All the frame vectors produced during preprocessing are stored in ImageDB. The Planner computes the CLIP embedding vector for the user query, and retrieves the top-$F$ frame vectors that are most similar to the user query vector. The frames corresponding to the top-$F$ vectors are the top-$F$ frames.
Context from the Planner includes two parts: Top-$N$ context chunks from TextDB, and the top-$F$ context frames from the ImageDB. As we show next, the Indexer does further processing to significantly reduce this context to at most $k$ context clips.
Figure 6: Indexer workflow
III-E Indexer
The Indexer refines the context from the Planner, and ensures that we have at most $k$ context clips for the Extractor.
@@ III-E Indexer
As stated earlier, $k$ is a configuration parameter of the iRAG system. A large $k$ implies a higher limit on the context clips, and the detailed extraction of so many clips will take time and adversely affect the quality of interactive querying. On the other hand, a very small $k$ can result in too few clips for detailed extraction, and the updated index may not have the context necessary for the LLM to provide a useful response to the user query. So, choice of $k$ is critical to ensure a balance between interactivity and usefulness of LLM responses.
Merge: As shown in Figure 6, the top-$N$ chunks from the Planner are further refined using a re-ranking algorithm (described next) to produce fewer context chunks. The merge component in the Indexer merges the top-$F$ frames and the clips corresponding to the context chunks by concatenation to produce at most $k$ context clips for the Extractor.
If the re-ranking algorithm is not used, then the clips corresponding to the top-$N$ chunks are merged with the top-$F$ frames to create at most $k$ context clips. However, if our new re-ranking algorithm is used, then a significant reduction in the top-$N$ clips is possible.
Re-ranking: A KNN classifier (described next) is used to filter the top-$N$ chunks into two classes: “accept” or “reject”. Figure 7 shows the re-ranking workflow. Each vector in the top-$N$ vectors is concatenated with the query vector, and the KNN classifier assigns an “accept” or “reject” label to the concatenated vector. Any top-$N$ vector that has the “reject” label is eliminated from further consideration. The remaining vectors form a new ordered list. The chunks corresponding to these vectors also form an ordered list, which is provided as input to the merge block.
@@
Training of KNN classifier: We train a classifier by using data from our baseline system, and iRAG’s Planner. Given a query, and a long video, we analyze the context retrieved by the baseline system, and iRAG. If a chunk retrieved by iRAG matches a chunk retrieved by the baseline system, then we assign a label of “1” (accept) to the iRAG chunk. Otherwise, we assign a label of “0” (reject) to the iRAG chunk. This creates labeled data that associates the concatenation of a user query vector and an iRAG chunk vector with a label “1” or label “0”. By using the test queries in VQA-2 dataset, we created labeled training data that was used to train our KNN classifier offline. After training, our KNN classifier assigns a label “1” to an unlabeled iRAG context chunk if 3 out of 5 neighbors of the concatenated query-chunk vector also have the label “1”.
Figure 7: Re-ranking workflow
III-F Extractor
The Extractor receives the context clips from the Indexer. It also receives (from the Planner) a list of AI models to run on each clip. The text data obtained after detailed extraction of the context clips is used to update TextDB. Then, the user query is processed again to create a prompt for the LLM. Figure 8 shows how the Planner, Indexer and Extractor work together in the iRAG system to incrementally update TextDB.
Figure 8: Interaction of key components in iRAG
IV Experimental results

### v2

# iRAG: Advancing RAG for Videos with an Incremental Approach

@@ Abstract.
Retrieval-augmented generation (RAG) systems combine the strengths of language generation and information retrieval to power many real-world applications like chatbots. Use of RAG for understanding of videos is appealing but there are two critical limitations. One-time, upfront conversion of all content in large corpus of videos into text descriptions entails high processing times. Also, not all information in the rich video data is typically captured in the text descriptions. Since user queries are not known apriori, developing a system for video to text conversion and interactive querying of video data is challenging.
@@
To address these limitations, we propose an incremental RAG system called iRAG, which augments RAG with a novel incremental workflow to enable interactive querying of a large corpus of videos. Unlike traditional RAG, iRAG quickly indexes large repositories of videos, and in the incremental workflow, it uses the index to opportunistically extract more details from select portions of the videos to retrieve context relevant to an interactive user query. Such an incremental workflow avoids long video to text conversion times, and overcomes information loss issues due to conversion of video to text, by doing on-demand query-specific extraction of details in video data. This ensures high quality of responses to interactive user queries that are often not known apriori. To the best of our knowledge, iRAG is the first system to augment RAG with an incremental workflow to support efficient interactive querying of a large corpus of videos. Experimental results on real-world datasets demonstrate 23x to 25x faster video to text ingestion, while ensuring that latency and quality of responses to interactive user queries is comparable to responses from a traditional RAG where all video data is converted to text upfront before any user querying.
@@
Generative AI, Retrieval Augmented Generation (RAG), Video Analytics, Large Language Models (LLMs), Vision Language Models (VLMs)
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3680088isbn: 979-8-4007-0436-9/24/10ccs: Information systems Information retrievalccs: Computing methodologies Natural language processing
1. Introduction
@@
Many applications routinely collect and store videos for offline analysis. For example, surveillance systems ensure public safety by monitoring public spaces, airports, transportation hubs, and critical infrastructure using video cameras (Shibata et al., 2023). These systems also store video feeds for offline analysis and investigations. Additionally, there is increasing usage of video analytics for patient monitoring in hospitals and healthcare facilities. Understanding information within videos is also essential for tasks like traffic monitoring, congestion management, and incident detection using feeds from traffic cameras (Mittal et al., 2023).
@@
Large language models (LLMs) like ChatGPT (Bubeck et al., 2023) have demonstrated their proficiency in human-like text-based dialogues. Inspired by the success of LLMs in understanding extensive text repositories (Arefeen et al., 2024a), recent proposals aim to extend this capability to video content analysis using LLMs (Showlab, 2023; Lin et al., 2023). These approaches employ vision AI models to analyze each video clip, outputting the visual information as text descriptions. For instance, an object detector AI model can identify objects such as cars, trucks, bicycles, and persons in each frame, and also specify their locations in human-readable text. These AI models can process lengthy videos, capturing visual information as a text document that concatenates descriptions from each clip. LLMs can then utilize these text documents through advanced techniques like Retrieval-augmented Generation (RAG) (Lewis et al., 2020), enabling them to provide insightful responses to interactive queries about the video content.
@@
Prior proposals to understand video content through interactive queries have two major limitations. First, the time required to process entire long videos and generate text descriptions by using heavyweight or complex AI models (which contain a large number of parameters and incur slow inference time) can be prohibitive. For example, it could take more than a day for heavyweight AI models to analyze visual information in a 24-hour surveillance video and generate text descriptions. This prevents the police from timely analysis of criminal incidents in the video. Second, not all visual information in the videos is captured in the text descriptions, even if we use state of the art heavyweight vision AI models (Carion et al., 2020; Taigman et al., 2014; Li et al., 2023c; Nguyen et al., 2022). Furthermore, the user queries are not known apriori, and it is difficult to determine which AI models should be used to convert visual content into text.
@@
To address the critical limitations of prior efforts to interactively query content in long videos, we propose iRAG. Unlike prior proposals (Showlab, 2023; Lin et al., 2023) that employ many AI models to extract textual information upfront, iRAG uses lightweight AI models (with fewer parameters and faster inference time) to quickly prepare an index for the content in the long videos. During preprocessing of the long video, our primary goal is to extract enough information from the video to be able to index the video data. Such an index facilitates effective extraction and retrieval of more detailed information from select sections of the video in response to interactive user queries.
@@
While utilizing lightweight AI models reduces the indexing time, the text descriptions may not be as rich, and many user queries may not be answerable by an LLM due to lack of facts in the text descriptions. To tackle this challenge, iRAG employs an incremental workflow. When the information extracted during indexing is inadequate to successfully respond to a user query, additional AI models are used on-demand to extract more details from selective parts of the long video. This strategy ensures that iRAG extracts necessary detailed textual information on-demand, addressing specific queries as they arise, rather than collecting all textual information upfront without knowing the specific details needed to answer arbitrary user queries.
iRAG has been deployed to process and analyze footage from the city-wide surveillance cameras. Since its deployment approximately one year ago, iRAG has significantly enhanced the capabilities of law enforcement agencies. By providing detailed insights, it has played a crucial role in investigating incidents, identifying suspects, and preventing potential crimes. iRAG’s ability to swiftly and accurately analyze vast amounts of video data has made it an invaluable tool in maintaining public safety and security.
@@
  1. (1)
    We introduce a novel incremental RAG system called iRAG, which enables immediate interactive querying of large video repositories. Unlike traditional methods that require many AI models upfront in a lengthy preprocessing step to convert video into text before any user query can be answered, iRAG quickly indexes large videos using lightweight AI models and employs select heavyweight AI models in a query-aware manner whenever needed.
  2. (2)
    We propose a Planner in the iRAG system that leverages indexed information and the user query to quickly identify query-specific heavyweight AI models and the relevant sections of the video for detailed analysis. We also propose an Extractor that uses these AI models to convert information from the relevant video sections into text descriptions and update the index. Incremental extraction of content from video is useful for eliciting responses from an LLM.
  3. (3)
    We evaluate iRAG on real-world datasets and demonstrate its capability to accelerate the video-to-text conversion process by a factor of 23-25x. Our novel incremental flow ensures that interactive user queries yield responses comparable in both latency and quality to those from a baseline traditional RAG system, where all video data is converted to text upfront before any user querying.
2. Incremental RAG
2.1. Prior work to leverage context in video
Prior work (Showlab, 2023; Lin et al., 2023) first produces a textual description of information in the entire video. The long video is divided into smaller clips. One or more generative AI models analyze these clips and produce textual descriptions for each. Concatenating these descriptions results in a long document. The entire preprocessing workflow is illustrated in Figure 1.
Figure 1. Video to text conversion
@@ 2.1. Prior work to leverage context in video
RAG: In Figure 1, a traditional Retrieval-augmented Generation (RAG) (Lewis et al., 2020) system is shown that augments the user prompt with query-specific context in the long document. The augmented prompt guides the LLM to produce more useful responses. We briefly review this process. The long text document is chopped up into chunks where a chunk can be a paragraph, or a sentence or some pre-determined number of consecutive characters. Each chunk is embedded into a vector of pre-determined length by an embedding algorithm, and these vectors are all stored in a vector database. All this makes it easy to identify chunks in the document that are similar to a user query. For instance, utilizing the embedding model, we can embed a user query into a vector. Then, by computing the dot product of the query vector with each chunk vector in the database, top-$k$ chunks that are most similar to the user query are retrieved to form a query-specific context from the long document utilizing the semantic search block in Figure 2.
Figure 2. Conventional RAG workflow.
Finally, the user query is augmented with query-specific context, and the augmented prompt results in a more contextually accurate response from an LLM.
2.1.1. Limitations of Prior work
@@
Long preprocessing time: Since user queries are not known apriori, many vision AI models are used to extract information from stored videos. However, models that capture more details are often computationally expensive. Preprocessing time is crucial, especially for time-sensitive tasks like crime analysis, where timely access to surveillance footage increases the chances of capturing critical evidence. As shown later, preprocessing a 24-hour surveillance video can take more than a day, delaying the use of this context in solving real-world problems and hindering law enforcement from gathering actionable intelligence, identifying suspects, and ensuring public safety.
@@
Figure 3. Time to generate a long document from real-world videos
Figure 3 shows the preprocessing bottleneck for querying and analyzing long videos from several New York City traffic intersections (Piadyk et al., 2023). We analyzed videos of 1, 7, and 30 days using the DETR object detection model (Carion et al., 2020) and the GRiT dense captioning model (Nguyen et al., 2022). On a server with an AMD Ryzen 5950X and NVIDIA GeForce RTX 3090, the average per-frame latency is about 70 ms for DETR and 1500 ms for GRiT. Processing one frame per second, the video-to-text generation for 1-day, 7-day, and 30-day videos takes approximately 1.57 days, 11 days, and 46.4 days, respectively. These lengthy conversion times hinder the development of practical systems for quickly querying and understanding long videos in real-world applications.
@@
2.1.2. Our new approach
Instead of executing all vision AI models upfront to extract textual information from entire videos, we propose an incremental RAG (iRAG) system that does query-aware, on-demand extraction of text data from select portions of the videos. Such a capability is not present in prior methods (Showlab, 2023; Lin et al., 2023) that solely rely on long preprocessing times to convert video to text once before any interactive querying of the long videos.
2.2. Preprocessing in iRAG
Unlike prior approaches, the goal of preprocessing in iRAG is to quickly build an index for the long video. For example, to answer the query “Is there a FedEx truck in a long video?”, the index (i.e. the object detection information) can be used to quickly identify the relevant clips where a truck was detected. Subsequent detailed extraction of text information from these selected clips (by using an AI model that can pick up text in images) can confirm if and where a FedEX truck is seen in the long video. We remark that a variety of indexes can be quickly constructed by using lightweight vision AI models such as DETR (Carion et al., 2020)-an object detection model and CLIP (Radford et al., 2021), which can retrieve clips based on user query. The heavyweight vision AI models such as GRiT (Nguyen et al., 2022) should be employed selectively, focusing only on video clips that align with the user’s specified query. This targeted incremental approach ensures optimal use of computational resources while maintaining high accuracy in content analysis.
2.3. Query-Response
Prompt 0.
@@ 2.3. Query-Response
Given a user query, iRAG retrieves the relevant context from the information collected during preprocessing by using the conventional RAG approach. Then, we construct a suitable prompt for the LLM by using the prompt template Prompt 1. iRAG invokes the incremental flow when the LLM is unable to answer a user query. Next, we describe the iRAG system, and its key components.
2.4. iRAG Components
Figure 4 provides an overview of the proposed iRAG system. The incremental workflow, which is activated only when the LLM cannot answer a query, is depicted within the light blue rectangle. There are two key components that support the query-aware, on-demand incremental flow: planner, and extractor.
Figure 4. iRAG overview: The light blue rectangle denotes additional workflow compared to a conventional RAG.
The iRAG Planner retrieves query-related clips by leveraging indexed information and employs AI models via the Extractor to gather additional information needed for generating a response to the user query. This incremental process can loop up to a preset maximum, although typically, one iteration suffices to answer most queries. The query response includes a textual answer along with supporting video clips, aiding users in verifying the accuracy of the LLM’s response. Future efforts may focus on integrating automatic response verifiers to ensure alignment between the textual response and video clips.
2.4.1. iRAG Planner
@@
The workflow in our planner is demonstrated in Figure 4. Given a query, the Planner computes a query embedding vector by using the same embedding algorithm that the RAG system used to embed preprocessing text chunks into vectors in TextDB (generated from the output of the DETR model (Carion et al., 2020)), and retrieves the top-$N$ chunks in TextDB that are most similar to the query embedding vector.
@@
Improved context retrieval: To enhance the quality of retrieved context, iRAG can maintain additional vector databases. For instance, iRAG can utilize ImageDB to store ”frame vectors” related to clips. Lightweight AI models like CLIP (Radford et al., 2021) embed visual information in each frame of the long video into a CLIP embedding vector, which serves as the frame vector. All frame vectors generated during preprocessing are stored in ImageDB. The Planner computes the CLIP embedding vector for the user query and retrieves the Top-$F$ frame vectors that are most similar to the user query vector. The frames corresponding to the Top-$F$ vectors are considered the Top-$F$ frames. The context from the Planner comprises two components: the Top-$N$ context chunks from TextDB and the Top-$F$ context frames from ImageDB. iRAG then identifies related clips associated with these contexts. The iRAG Planner conducts further processing to reduce the top-$N$ chunks to a maximum of $k$ chunks, ensuring prompt and interactive query responses to the user.
@@
Context Chunks Filtering. After retrieving the context chunks, there may still be irrelevant clip information included. Extracting irrelevant video clip information at the incremental stage will add latency to the response. Therefore, the Planner ensures that we have at most $k$ important context clips associated to the chunks for the Extractor. A large $k$ implies a higher limit on the context clips, and the detailed extraction of so many clips will take time and adversely affect the quality of interactive querying. On the other hand, a very small $k$ can result in too few clips for detailed extraction, and the updated index may not have the context necessary for the LLM to provide a useful response to the user query. So, choice of $k$ is critical to ensure a balance between interactivity and usefulness of LLM responses.
Classifier: iRAG employs a KNN classifier to group the top-$N$ chunks into two classes: “accept” or “reject”. The embedding of each top-$N$ check is concatenated with the query vector to form the feature vector for the classifier, which assigns an “accept” or “reject” label to each concatenated vector. Any vector having “reject” label is eliminated from further consideration. From the accepted chunks, the top $k$ chunks most similar to the query are selected for extraction.
@@
Training of KNN classifier: We train a classifier by using data from our baseline system, and iRAG’s Planner. Given a query, and a long video, we analyze the context retrieved by the baseline system, and iRAG. If a chunk retrieved by iRAG matches a chunk retrieved by the baseline system, then we assign a label of “1” (accept); otherwise, we assign a label of “0” (reject). Using the test queries in the VQA-2 dataset (vqa, 2024), we created labeled training data for our KNN classifier offline. After training, our KNN classifier assigns a label “1” to an unlabeled iRAG context chunk if 3 out of 5 neighbors of the concatenated query-chunk vector have the label “1”.
2.4.2. iRAG Extractor
The Extractor receives context clips from the Planner, along with a list of heavyweight AI models to run on each clip. Instead of converting the entire video to text using a heavyweight model such as GRiT (Nguyen et al., 2022), the Extractor runs the heavyweight model only on the selected context clips deemed essential by the Planner for accurate query responses. The detailed text data extracted from these context clips is then used to update the TextDB vector database. Subsequently, the user query is re-issued with the updated context to generate a prompt for the LLM to obtain the response. Figure 4 illustrates how the Planner and Extractor collaborate in the iRAG system to incrementally update the TextDB.
3. Experimental results